OpenStudy (anonymous):

What is the volume of a triangular prism that has a height of 9 meters, length of 14 meters, and base of 4 meters?

OpenStudy (lgbasallote):

 \[V = \frac{bhl}{2}\] where: V = volume b = base h = height l = length does that help?

OpenStudy (anonymous):

I appreciate your help, but I'd personally like a straight answer on how to solve it, not just a drawing.

OpenStudy (anonymous):

just follow the formula..

OpenStudy (lgbasallote):

i gave you how to solve it

OpenStudy (lgbasallote):

i gave you a drawing AND a formula

OpenStudy (anonymous):

Never mind. Thanks anyway.

OpenStudy (anonymous):

\[V = \frac{ 4 m \times 14 m \times 9 m }{ 2 } = 252 m\]

Directrix (directrix):

V = B*h where B is the area of the triangular base and h is the height of the prism. The area the triangular base is 1/2 * 4 * 9 = 18 square meters The volume of the prism is: V = (18)(14) = 252 cubic meters. Volume is expressed in cubic units.

OpenStudy (anonymous):

haha... i'm forget the cubic meters.. thank you Directrix